8. A ball rolls horizontally with a speed 0f 7.6 m/s off the edge of a tall platform. If the ball lands 8.7 m from the point on the ground directly below the edge of the platform, what is the height of the platform?



9. A stone thrown off a bridge 20 m above a river has an initial velocity of 12 m/s at an angle of 45° above the horizontal.


What is the range of the stone?


At what velocity does the stone strike the water?



1
Expert's answer
2022-03-15T16:50:08-0400

Solution.

"8. v=7.6 m\/s;"

"L=8.7m;"

"L=vt\implies t=\dfrac{L}{v};"

"t=\dfrac{8.7}{7.6}=1.44s;"

"h=\dfrac{gt^2}{2};"

"h=\dfrac{9.8\sdot1.44^2}{2}=10.16m;"

"9.y_0=20m;"

"v=12 m\/s;"

"\alpha=45^o;"

"y=y_o+v_{0y}t-\dfrac{gt^2}{2};"

"0=20+12\sdot sin45^ot-\dfrac{9.8\sdot t^2}{2};"

"t=3.06s;"

"L=v_{ox}t;"

"L=12\sdot0.7071\sdot3.06=26m;"

"v_{y}=v_{0y}-gt;"

"v_y=12\sdot0.7071-9.8\sdot3.06=-21.5m/s;"

Answer: "8.h=10.16m;"

"9. L=26m; v_y=-21.5 m\/s."